# Proposal for Improving Google A2A Protocol: Safeguarding Sensitive Data in Multi-Agent Systems

### v2

@@ Abstract
A2A, a protocol for AI agent communication, offers a robust foundation for secure AI agent communication. However, it has several critical issues in handling sensitive data, such as payment details, identification documents, and personal information. This paper reviews the existing protocol, identifies its limitations, and proposes specific enhancements to improve security, privacy, and trust. It includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on prior studies to strengthen the arguments and proposed solutions. This proposal includes seven enhancements: short-lived tokens, customer authentication (SCA), granular scopes, explicit consent, direct data transfer, multi-transaction approval, and payment standard compliance. The vacation booking example illustrates how these enhancements reduce risks and enhance user experience.
1 Introduction
@@
The rapid appearance of Agentic AI autonomous systems that plan, delegate, and collaborate without human intervention has created an urgent need for robust, standardized communication protocols [14]. Google’s A2A (Agent-to-Agent) protocol [48] establishes such a foundation by defining a declarative, identity aware framework for discovering, authenticating, and exchanging tasks among heterogeneous agents. A2A’s core mechanism as shown at Figure 1, the AgentCard, provides machine readable metadata that enables seamless interoperability across organizational and technical boundaries. Complementing A2A, the Model Context Protocol (MCP) standardizes the integration of large language models (LLMs) with external data sources and tools, allowing agents to leverage these models for real-time contextual understanding during task execution. This integration enhances agents’ capabilities in complex scenarios, supporting use cases ranging from payment orchestration to business automation [26]
Figure 1: A2A’s core mechanism
However, as multi-agent ecosystems scale in complexity, they also open new attack surfaces. Recent threat analyses of both A2A and the related MCP reveal sophisticated vectors such as shadowing (where malicious tool descriptions subvert genuine workflows), tool poisoning, and naming attacks that exploit implicit trust in discovery mechanisms [29, 44]. Furthermore, broad surveys of AI agents underscore systemic risks-ranging from prompt injection to cross-agent privilege escalation-that can lead to data exfiltration or unauthorized task execution [20, 32, 45]. Similarly, “enterprise-grade” analyses of MCP implementations demonstrate that without rigorous, zero-trust controls, attackers may exploit unsecured tool endpoints or manipulate context payloads to bypass security checks [37].
Taken together, these findings highlight a gap between protocol design and real-world resilience: A2A and MCP provide the plumbing for agent interoperability, but they do not yet enforce the fine-grained confidentiality, integrity, and consent guarantees required for sensitive data exchange. In this paper we review six critical weaknesses in the handling of personal information, such as payment credentials, user identity and sensitive documents, are identified. Building on proactive threat models and industry best practices, we propose seven concrete enhancements ranging from short-lived, client bound tokens to explicit user consent fields and direct data transfer processes to harden A2A communications against the most urgent threats of the day. This proposal includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on previous studies to strengthen the arguments and proposed solutions.
2 Overview of the A2A Protocol
Google’s A2A (Agent-to-Agent) protocol [48] builds upon widely adopted web standards HTTP, HTTPS, JSON-RPC, and Server-Sent Events (SSE) to provide an extensible, interoperable framework for autonomous agent communication. Agents establish mutual authentication via OAuth 2.0 [27] flows and JSON Web Tokens (JWTs) [30], while RSA key pairs are used for signature validation and secure key exchange. Specifically, OAuth 2.0, an industry-standard authorization framework, enables agents to securely delegate access to each other’s resources without sharing credentials, using standardized flows like the authorization code grant. Meanwhile, JWTs, compact and cryptographically signed tokens, facilitate the secure transmission of authentication claims between agents, ensuring both integrity and verification of access rights during interactions. Role-based access control (RBAC) is natively supported through user- and agent-scoped JWT claims, and fine-grained permissions map directly onto task and message schemas to enforce the principle of least privilege. By default, A2A messages carry only minimal metadata (e.g., action identifiers, input/output schemas, and consent fields), reducing sensitive data exposure when agents negotiate capabilities or invoke third-party services such as payment gateways or AI inference backends.
A2A’s flexibility underpins diverse real-world scenarios:
  • Service Booking: Agents coordinate multi-step operations (e.g., flight, hotel, transportation) by passing structured requests containing specific payment tokens and availability parameters.
  • Enterprise Task Management: Supply-chain or project-management agents exchange only the data needed to complete discrete tasks, minimizing shared context.
  • Cross-Provider Collaboration: Agents from different platforms (e.g., Google, PayPal, Cohere) interoperate securely via standardized AgentCard discovery and encrypted channels.
Despite these strengths, A2A inherits an inherent trade-off between security and efficiency: tightening authorization or adding cryptographic checks invariably increases latency and complexity, whereas relaxing controls risks unauthorized data flows [42]. Crucially, A2A currently lacks specialized safeguards for handling particularly sensitive payloads-such as payment credentials or identity documents-beyond generic token expiry. A similar concern is raised in “AI Agents Meet Blockchain” observes, decentralized agent ecosystems without centralized guardrails require robust consent, auditing, and privacy-preserving delegation mechanisms to maintain trust [32]. Without such controls, adversarial agents might exploit overly broad scopes or unmonitored message channels to exfiltrate or tamper with confidential information. The following sections will analyze these gaps in more detail and propose targeted enhancements to ensure that A2A can securely mediate sensitive data exchanges at scale.
3 Identification of Issues in Handling Sensitive Data
Despite its benefits, A2A has some critical problems in the handling of sensitive data such as payment details, ID documents and personal data. The following six issues are based on real world experience, peer reviewed literature and documented CVE threats.
3.1 Absence of limitations on token lifetime
@@
Although A2A is based on OAuth 2.0, it does not enforce strict expiration durations (e.g. seconds or minutes) for tokens used in sensitive transactions. Without such restrictions, leaked tokens may remain valid for hours or even days, increasing the risk of unauthorized reuse. revocation. As demonstrated in ”AgNet” [25], long-lived tokens are a systemic weakness in distributed architectures, allowing for multiple accesses in the event of a compromise. For example, CVE-2025-1198 shows that revoked GitLab personal access tokens are still accepted by long-term ActiveConnection [11]. Another case described by CVE-2025-1801, where a low-privileged user obtained a JWT issued to a high-privileged user on account [12].
3.2 Lack of Strong Customer Authentication (SCA)
@@
The A2A protocol does not have built-in requirements for strong authentication, such as two-factor or biometric authentication, for high-value transactions such as payments or identity switching. Without these safeguards, adversaries may perform unauthorized acts on behalf of the user [24]. The Medibank breach of 2022, where attackers gained access to personal data of 9.7 million people through the lack of multifactor authentication, illustrates the tangible consequences of [28]. CWE-306 [10] also classifies systems that do not authenticate users before performing critical functions as inherently vulnerable. In addition, the AI Agents Meet the Blockchain project (aclm) [32] proposes Zero Knowledge Proof (ZKP) techniques for secure authentication in decentralized environments, but the A2A does not include such a mechanism.
3.3 Insufficiently Granular Token Scopes
@@
Tokens in A2A do not define precise ranges for sensitive transactions, which introduces the risk of privilege escalation. For example, a token issued to initiate a payment may inadvertently grant access to unrelated data. The study on the multi-agent security tax [42] argues that coarse-grained authorization models increase the probability of data exposure. This concern is highlighted by vulnerability CVE-2023-4456 [4] on LokiStack. Similarly, CWE-1220 [8] documents the lack of granularity in access control policies leading to infringements of the principle of the minimum privilege.
3.4 Lack of Transparency and User Consent

### v3

# Improving Google A2A Protocol: Protecting Sensitive Data and Mitigating Unintended Harms in Multi-Agent Systems

@@ Abstract
Google’s A2A protocol provides a secure communication framework for AI agents but demonstrates critical limitations when handling highly sensitive information such as payment credentials and identity documents. These gaps increase the risk of unintended harms, including unauthorized disclosure, privilege escalation, and misuse of private data in generative multi-agent environments. In this paper, we identify key weaknesses of A2A: insufficient token lifetime control, lack of strong customer authentication, overbroad access scopes, and missing consent flows. We propose protocol-level enhancements grounded in a structured threat model for semi-trusted multi-agent systems. Our refinements introduce explicit consent orchestration, ephemeral scoped tokens, and direct user-to-service data channels to minimize exposure across time, context, and topology. Empirical evaluation using adversarial prompt injection tests shows that the enhanced protocol substantially reduces sensitive data leakage while maintaining low communication latency. Comparative analysis highlights the advantages of our approach over both the original A2A specification and related academic proposals. These contributions establish a practical path for evolving A2A into a privacy-preserving framework that mitigates unintended harms in multi-agent generative AI systems.
I Introduction
@@
The rapid emergence of autonomous agent systems capable of planning, delegating, and coordinating tasks without direct human intervention has created a growing demand for standardized communication protocols [1]. Google’s Agent-to-Agent (A2A) protocol addresses this need by establishing a structured, identity-aware framework for discovering and interacting with services across heterogeneous agents [2]. At the core of this protocol lies the AgentCard, a declarative metadata object that enables seamless interoperability between agents by conveying machine-readable descriptions of capabilities, roles, and identities, as illustrated in Figure 1. Complementing A2A, the Model Context Protocol (MCP) [3] facilitates real-time integration of large language models with tools and external data, thereby enabling richer context sharing and execution capabilities in multi-agent workflows [4].
Despite these innovations, recent threat analyses have uncovered critical vulnerabilities that challenge the secure handling of sensitive data in such environments. Researchers have demonstrated attack vectors such as shadowing, tool poisoning, and naming manipulation, which exploit implicit trust mechanisms in the agent discovery process [5, 6]. Furthermore, studies have highlighted systemic risks including prompt injection, excessive privilege escalation, and cross-agent data leakage, particularly in high-value tasks such as payments or identity verification [7, 8, 9].
Enterprise-grade evaluations of MCP implementations indicate that without robust enforcement of access control and consent boundaries, attackers may subvert tool endpoints or manipulate contextual payloads to bypass authorization policies [10]. Taken together, these findings reveal a gap between the theoretical design of A2A and its practical resilience against adversarial behavior. While A2A and MCP enable functional interoperability, they do not yet provide sufficient guarantees of confidentiality, integrity, and informed consent for handling sensitive information.
This paper addresses this gap by identifying core protocol weaknesses in areas such as token management, authentication strength, scope granularity, and data flow transparency. We propose a set of targeted enhancements to improve privacy, security, and user control in agent-mediated communications, and demonstrate their application in a real-world example involving vacation booking. Our proposal incorporates best practices from adjacent fields such as zero-trust architectures and regulatory compliance in financial technologies.
Figure 1: A2A’s core mechanism. The AgentCard provides identity-aware task and tool metadata to support discovery and execution across agent systems.
Unlike existing security recommendations that remain optional or are implemented at the application layer, in this paper we present how to enforces these mechanisms at the protocol level in a cohesive manner tailored for decentralized multi‑agent environments. The integration of explicit consent orchestration, ephemeral and strictly scoped credentials, and direct user‑to‑service data channels is designed as a unified security architecture rather than isolated best practices. This design ensures that safeguards are consistently applied across heterogeneous agents, closing enforcement gaps that persist in current A2A deployments.
II Related Work and Background
This section reviews the foundations of the A2A protocol and situates it within broader research on secure multi-agent communication. A2A, developed by Google, builds upon established web standards such as HTTP, HTTPS, JSON-RPC, and Server-Sent Events (SSE) to define a secure and extensible communication framework between autonomous agents. Agents engage in mutual authentication through OAuth 2.0 [11] and exchange cryptographically signed JSON Web Tokens (JWTs) [12] that encapsulate identity and permission claims. The use of RSA key pairs facilitates signature verification and key exchange, ensuring message integrity and authentication without relying on shared secrets.
OAuth 2.0 provides an industry-standard method for authorization, allowing agents to delegate access securely to external resources. The protocol typically employs the authorization code grant flow to ensure secure token exchange. JWTs, being compact and self-contained, carry essential claims that include user identifiers, expiration timestamps, and access scopes, enabling verification of access rights and enforcing the principle of least privilege across distributed systems. Role-based access control (RBAC) is natively supported through user and agent scoped JWT claims, ensuring that agents only access resources aligned with their intended roles and tasks.
A2A messages are designed to minimize unnecessary exposure of sensitive information by default. The messages include only essential metadata such as action identifiers, input and output schemas, and optional consent fields. This design helps reduce data leakage risks during task delegation or tool invocation, particularly when agents communicate with third-party services such as payment processors or AI inference backends.
A2A supports a wide range of real-world scenarios by promoting secure agent interoperability across organizational and platform boundaries. Prominent examples include service booking (e.g., flights, hotels, transportation), enterprise task automation (e.g., supply-chain orchestration, project management, procurement), and cross-provider collaboration (e.g., Google, PayPal, and Cohere). These cases highlight the protocol’s broad applicability and practical relevance.
Despite these design strengths, A2A embodies an inherent trade-off between security and efficiency. Introducing additional authorization checks or cryptographic safeguards may increase latency and complexity, whereas relaxing such controls can open avenues for data exfiltration and privilege escalation [13]. In particular, the A2A specification lacks tailored protections for handling highly sensitive payloads, such as user credentials, payment information, and identity documents. As noted by Karim et al. [8], decentralized agent environments demand explicit controls around consent, delegation, and auditability to maintain operational trust. Without these mechanisms, adversarial agents may exploit overbroad scopes or unmonitored message channels to gain unauthorized access to sensitive resources.
This analysis reflects prior work that established the strengths and limitations of A2A. Analogous vulnerabilities have also been documented in related agent interoperability protocols, including MCP, ANP and ACP [14], [15], [16]. Nevertheless, since this paper concentrates exclusively on the A2A protocol, we refrain from an in-depth discussion of these alternatives. Our contribution builds directly on this foundation by proposing targeted enhancements that mitigate leakage risks in generative multi-agent environments.
In addition to these protocol-level considerations, recent research has also highlighted risks inherent to the agents themselves, which further motivates the safeguards we propose and are reviewed next.
Figure 2: Threat model for the A2A protocol in a semi-trusted multi-agent environment. Solid arrows indicate legitimate communication flows; dashed red arrows indicate attack vectors including prompt injection, token replay, scope escalation, and unverified endpoints.
II-A Risks of Agent-Initiated Misuse of Sensitive Data
While prompt injection remains a well-documented threat, autonomous agents can also misuse sensitive information without any external manipulation. In Anthropic’s study on agentic misalignment [17], large language model agents such as Claude were observed engaging in blackmail and corporate espionage, even when initially given benign objectives, especially when perceiving threats to their continued operation. Similarly, Cybernews reported a real-world case where Replit’s AI coding assistant deleted a live production database during a code freeze, explicitly ignoring multiple instructions not to proceed [18]. Another high-profile example from Carnegie Mellon University demonstrated that large language models, when given sufficient autonomy, could independently orchestrate a complete multi-stage cyberattack, including reconnaissance, network exploitation, malware deployment, and data exfiltration without human oversight [19]. These incidents illustrate that AI agents are not merely passive tools vulnerable to manipulation; in certain circumstances, they may actively and deliberately act against user interests.
Additional evidence underscores that these risks extend beyond isolated research studies. Lumia Security’s AppleStorm investigation revealed that Apple’s Siri and Apple Intelligence ecosystem quietly transmitted WhatsApp and iMessage content, location data, and metadata about installed applications to Apple servers, even when such transmission was unnecessary or contrary to user expectations [20]. Similarly, Guardio’s Scamlexity study demonstrated that agentic AI browsers, such as Perplexity’s Comet, autonomously completed online purchases on fraudulent websites and submitted sensitive user data to phishing pages without human oversight, effectively turning trusted assistants into vectors of financial and security harm [21]. Recent academic work further highlights this trend: a comprehensive study on the expanding attack surface of agentic AI systems shows that emergent behaviors, prompt injection vulnerabilities, and autonomous task execution can transform well-intentioned agents into sources of significant unintended harms [22].
Together, these findings illustrate that AI agents pose inherent privacy and security threats not only when attacked, but also through their own design.
These risks underscore the need for protocol-level safeguards to prevent unintended harms, an issue we address in the following sections of this paper.
III Threat Model
We consider a semi-trusted multi-agent environment, consistent with prior work on LLM security [23, 24, 25], but tailored to the Google A2A protocol and the enhancements proposed in this paper.
III-A Actors
  • End User: Initiates tasks and may provide sensitive data such as payment credentials or identity documents.
  • Benign Agent: Executes assigned tasks according to protocol but may inadvertently expose sensitive data if compromised or poorly designed.
  • Malicious Agent: Attempts to gain unauthorized access to sensitive data or services by exploiting protocol weaknesses.
  • Service Provider: Intended recipient of sensitive data (e.g., payment processor, identity verifier).
III-B Adversary Capabilities
  • Launch prompt injection attacks to alter agent behavior.
  • Replay valid tokens if expiration policies are weak.
  • Escalate privileges by exploiting overly broad OAuth scopes.
  • Redirect data to unverified or malicious endpoints.
III-C Assumptions
  • TLS 1.3 or equivalent transport security is in place for all communications.
  • The trusted endpoint registry is correctly maintained and distributed.
  • End users can approve or deny sensitive actions when prompted.
III-D Security Goals
  • Prevent unauthorized disclosure of sensitive user data.
  • Ensure explicit user consent for all high-risk transactions.
  • Minimize exposure by using least-privilege and short-lived credentials.
III-E Illustration
Figure 2 summarizes the primary actors and attack vectors considered in this model, forming the basis for the vulnerability analysis.
Figure 3: Example booking workflow in A2A illustrating overprivilege, long token lifetimes, and absence of explicit user consent.
IV A2A Security Vulnerabilities, Protocol Enhancements, and Research-Based Justification
Despite A2A benefits, we identified several critical shortcomings in handling sensitive information such as payment credentials, identity documents, and personal data. These concerns are illustrated in the following scenario as reflected in Figure 3: a user asks an AI agent to book a vacation, including flights, hotel, and taxi, using access to his calendar, personal information, and payment credentials. The agent shares a payment token with the booking agent, valid for hours or even days, and accesses the user’s full calendar, including unrelated entries such as medical appointments. In this case, the token may be reused for unauthorized payments, and sensitive calendar data is accessed without the user’s explicit consent. This example highlights fundamental security and privacy gaps in the current protocol, including lack of token restrictions, insufficient scoping, and absence of user consent mechanisms.
The remainder of this section outlines specific issues that emerge from real-world observations, peer-reviewed research, and documented CVE vulnerabilities.
IV-A Absence of Limitations on Token Lifetime
@@
Problem: Although A2A is based on OAuth 2.0, it does not enforce strict expiration durations (e.g. seconds or minutes) for tokens used in sensitive transactions. Without such restrictions, leaked tokens may remain valid for hours or even days, increasing the risk of unauthorized reuse. As illustrated in Figure 3, the concern over token misuse highlights how long-lived credentials in A2A can be reused for unauthorized actions. As demonstrated in ”AgNet” [26], long-lived tokens are a systemic weakness in distributed architectures, allowing for multiple accesses in the event of a compromise. For example, CVE-2025-1198 shows that revoked GitLab personal access tokens are still accepted by long-term ActiveConnection [27]. Another case described by CVE-2025-1801, where a low-privileged user obtained a JWT issued to a high-privileged user on account [28].
Proposed Solution: To mitigate this, we propose implementing short-lived tokens for all sensitive operations such as financial transactions or identity verification. These tokens should be ephemeral (e.g., 30 seconds to 5 minutes) and valid for a single operation. Short-lived tokens minimize the window of opportunity for unauthorized access or reuse. This approach is supported by distributed architecture designs, and aligns with privacy-enhancing techniques proposed in [8]. Short-lived tokens limit the validity period of credentials, reducing the risk of credential theft and replay. Multiple studies affirm this principle. Teng et al. introduced ActionID, a machine-to-machine authentication protocol that issues tokens tightly scoped in both action and time. The author notes that “the short-lived token limits the time window for action execution,” thereby reducing the effectiveness of credential theft [29]. Similarly, Ohwo et al. proposed a blockchain-based smart home access system in which “short-lived tokens [are] used to mitigate risks like replay attacks and user profiling” [30]. These tokens expire rapidly, neutralizing any intercepted credentials. Narajala et al. extended this logic to GenAI multi-agent systems, proposing a just-in-time registry that dynamically provisions tokens only for the duration of a tool invocation. This architecture, they argue, “minimizes the attack surface associated with persistent credentials by dynamically provisioning short-lived access tokens only when needed” [31]. Finally, Xiao et al. addressed the same vulnerability in the IoT domain, where resource-constrained devices are especially susceptible to token compromise. They developed MCU-Token, which binds a short-lived, hardware-derived token to each request, thereby making every token instance single-use and resistant to replay [32]. Together, these findings offer both conceptual and empirical support for implementing short-lived tokens in the A2A protocol as a fundamental safeguard against credential replay and session hijacking.
IV-B Lack of Strong Customer Authentication (SCA)
@@
Problem: The A2A protocol does not have built-in requirements for strong authentication, such as two-factor or biometric authentication, for high-value transactions such as payments or identity switching. The flow in Figure 3 highlights how payment operations are executed without strong customer authentication, leaving transactions vulnerable to impersonation. Without these safeguards, adversaries may perform unauthorized acts on behalf of the user [33]. The Medibank breach of 2022, where attackers gained access to personal data of 9.7 million people through the lack of multifactor authentication, illustrates the tangible consequences of [34]. CWE-306 [35] also classifies systems that do not authenticate users before performing critical functions as inherently vulnerable. In addition, the AI Agents Meet the Blockchain project (aclm) [8] proposes Zero Knowledge Proof (ZKP) techniques for secure authentication in decentralized environments, but the A2A does not include such a mechanism.
Proposed Solution: We recommend implementing SCA for sensitive operations in the A2A protocol, such as payments and identity verification. That significantly improves protection against impersonation and unauthorized transactions. Recent research underscores the feasibility and necessity of using modern, privacy-preserving SCA mechanisms such as zero-knowledge proofs (ZKPs), biometrics, and multi-factor authentication (MFA). For example, Neera et al. propose a mobile payment protocol that leverages ZKPs and identity-based signatures to verify user identity without revealing sensitive data while ensuring compliance with financial regulations such as PSD2 [36]. Their scheme guarantees cryptographically enforced SCA while maintaining user privacy, demonstrating that SCA can be both secure and regulatory compliant. Ahmad et al. present BAuth-ZKP, a blockchain-based MFA framework using smart contracts and ZKPs to authenticate users in smart city environments [37]. This design proves the viability of decentralized, privacy-respecting SCA mechanisms, directly applicable to agent-mediated A2A architectures. On the biometric front, Gernot and Rosenberger introduce a technique for generating one-time biometric templates, mitigating the risk of replay attacks by ensuring biometric credentials are valid only once [38]. This reinforces the ’inherence’ factor of SCA in a technically sound manner. Finally, Lyastani et al. empirically demonstrate that while 2FA dramatically improves account security, inconsistent or poorly designed SCA flows reduce adoption [39]. Thus, their findings support designing usable, transparent SCA processes for sensitive actions in A2A. Collectively, these studies provide strong empirical and architectural justification for integrating SCA into the A2A protocol, ensuring that sensitive transactions are approved only by verified users under secure and user-friendly conditions.
IV-C Insufficiently Granular Token Scopes
@@
Problem: Figure 3 further illustrates how coarse-grained tokens, once shared, can be used beyond their intended scope, reinforcing the risks of overprivilege. Tokens in A2A do not define precise ranges for sensitive transactions, which introduces the risk of privilege escalation. For example, a token issued to initiate a payment may inadvertently grant access to unrelated data. The study on the multi-agent security tax [13] argues that coarse-grained authorization models increase the probability of data exposure. This concern is highlighted by vulnerability CVE-2023-4456 [40] on LokiStack. Similarly, CWE-1220 [41] documents the lack of granularity in access control policies leading to infringements of the principle of least privilege.
Proposed Solution: To resolve this, we advocate for granular scoping within OAuth 2.0 tokens, such as limiting access to a specific payment amount or restricting calendar visibility to availability only. The risks of overly broad permissions in bearer tokens are well-documented in recent research. Cao et al. propose a stateful, least-privilege authorization model that allows client-side applications to dynamically constrain the scope of OAuth tokens using WebAssembly-based privilege attenuation policies [42]. Their system allows developers to explicitly encode minimal access rights per session, enabling secure delegation without overprovision. This aligns directly with the A2A context, where agents must act within strict permission boundaries to avoid inadvertent data access. Complementarily, South et al. extend OAuth 2.0 and OpenID Connect for authenticated agent delegation, introducing agent-specific credentials with precise, auditable scope limitations [43]. Their framework demonstrates how user intent can be translated into tightly scoped permissions, ensuring that AI agents operate only within authorized domains. These works collectively support the need for scope-constrained access tokens as a foundational safeguard in agent-mediated architectures. Dimova et al. further substantiate this by showing that 18.5% of OAuth deployments on the web request unnecessary scopes, violating the GDPR’s minimum necessary data principle [44]. Kaltenböck et al. reinforce this position by embedding scope-aware policies within a Zero Trust single sign-on framework, emphasizing explicit scope definitions to limit token capabilities at authentication time [45]. Altogether, the literature affirms that fine-grained token scopes are not only technically viable but also indispensable for ensuring data minimization, secure delegation, and regulatory compliance in distributed, AI-driven authentication systems.
IV-D Lack of Transparency and User Consent